I am trying to help my 10 year old with some word problems, but they seem vague and i do not want to lead him astray.
"In 6 years Megan will be 3 times as old as David. David is 4 now. How old is Megan?"
Is it asking for Megan's age NOW or 6 years from now? This is what i have come up with before i explain it to him: D=David's Age, M=Megan's Age.
D+6=10 (david's age plus 6years from now) therefore M=(D+6)x3 Megan's age equals Davids age plus 6 years from now three times older so M=30? or is it as easy as, David is 4 and Megan is David times Three, D=4, M=D3(M=4x3), M=12.
The question is exactly off his worksheet. Any help would be appreciated.
bryan.walton@cox.net
Thanks,
Bryan

Answer by mananth(16946)

"In 6 years Megan will be 3 times as old as David. David is 4 now. How old is Megan?"

..
David = 4 years now.
after 6 years
David = 6+4 = 10 years
Megan = 3*10 = 30 years.
..
Megan's age now will be 30-6= 24 years.
...
m.ananth@hotmail.ca